tn Hawaii shark attacks sre very rare. Prthermore, bekR SS% of shark attacks ore won, ba not include brie, statistical profte regarding the most recent rine hark etsek. ofwMh wee nothtal. เฮ , be the ettacks (a) Make a histogram for the probability distribution of r 030 030 0 25 0.20 ois 010 25 015 005 0 25 0 20 015 25 0 20 0.15 ถเอ 0 05 005
(b) What is the probability all nine shark attacks are nonfatal? (Round your answer to three decimal places.) (c) What is the probability that seven or more of the nine shark attacks are nonfatal? (Round your answer to three decimal places) What is the expected number of nonfatal shark attacks out of the nine? What is the standard deviation of the r-probability distribution? (Round your answer to three decimal places.) Need Help? Read
Answer #1

Ans:

a)

P(x=k)=9Ck*0.55k*0.459-k

x P(x)
0 0.001
1 0.008
2 0.041
3 0.116
4 0.213
5 0.260
6 0.212
7 0.111
8 0.034
9 0.005

P(x) 0.300 0.250 0.200 0.150 0.100 0.050 0.000 0 1 2345 6 789

last option is correct.

b)

P(x=9)=0.55^9=0.005

c)

P(x>=7)=P(x=7)+P(x=8)+P(x=9)

=9C7*0.557*0.452+9C8*0.558*0.451+9C7*0.559*0.450

=0.111+0.034+0.005

=0.150

d)

Expectation=np=9*0.55=4.95

Standard deviation=sqrt(np(1-p))=sqrt(9*0.55*0.45)=1.492